For the first two lines of Paschen series, identify the value of n1 and n2.
Calculate the wave number values for these lines.
1
Expert's answer
2019-02-05T02:08:05-0500

In Paschen series the elctrons fall back to n1 = 3.

First line n1 = 3 , n2= 4

We should use Rydberg equation to calculate wave number value:

ν = R*(1/n12 - 1/ n22)

v = 1.09678*107 (1/32 - 1/42) = 5.331*105 m-1

Second line n1 = 3 , n2 = 5

v = R*(1/n12 - 1/ n22)

v = 1.09678*107 (1/32 - 1/52) = 7.799*105 m-1